3x+3y+3z=12
2x+2y+2z=9
x+y+z=6

Answer by Alan3354(69443)

3x+3y+3z=12
2x+2y+2z=9
x+y+z=6
----------
These equations are inconsistent.
Eqn 3 times 2 --> 2x+2y+2z= 12, but Eqn 2 says = 9
There is no solution.